# Predicting potentially unfair clauses in Chilean terms of services with natural language processing

### v1

@@ 6.3 Limitations and Outlook
We see several important directions for future work. First, open questions remain as to potential improvements when combining the most advanced architectures and training methodologies, such as DeBERTaV3, with a suitable language- and domain-specific training corpora, such as used by RoBERTalex or BETO. Analogously to domain-specific pre-training, [24] propose the Lawma model by fine-tuning Llama-3 8b on 260 legal tasks, effectively emulating pre-training. Their Lawma outperforms GPT-4 in the zero- and three-shot settings. We consider both Spanish legalese pre-training and task-specific fine-tuning of LLMs with more than a billion parameters an interesting avenue for future work, as [24] report increased performance from only a few hundred samples. The findings of [5] on the limited reasoning capabilities of LLMs even necessitate enhancements of today’s generative LLMs beyond zero-shot learning.
@@ 7 Conclusion
A growing number of products and services are offered through online platforms. However, their Terms of Service remain enigmatic to consumers. Even if they would read them, there is no alternative but to accept them. This information asymmetry and rational apathy present in modern consumer contracts may result in uncontested but potentially unfair clauses and their adverse effects on consumers. It follows that there is a growing interest in the potential of automatic contract analysis.
We present an novel annotation scheme and a substantial dataset comprising 50 online Terms of Service. The corpus encompasses two tasks for three categories of potentially unfair clauses. Our extensive evaluation of a diverse set of transformer-based methods reveals that multiple factors impact the detection and classification rates, including language- and domain-specific pre-training, the number samples in few-shot prompts, and architectural considerations. Potentially unfair clauses are detected with macro-F1 scores ranging from between $79\%$ to $89\%$ and micro-F1 scores of $95\%$ to $96\%$. The highest classification rates range from between $60\%$ to $70\%$ macro-F1 score and $64\%$ to $80\%$ micro-F1 score. This underscores the wide range of difficulties inherent in the novel datasets.
This work introduces the first Spanish-language classification dataset of legal clauses. Moreover, it applies law from outside a major jurisdictions, specifically the Chilean legal code. We also present the first comprehensive study on Spanish-language and multilingual transformer-based models on legal classification tasks in Spanish. This paper serves as the foundation for substantial future research and the implementation of its methodology to benefit Chilean and Latin American consumers.
Declarations
[…]
@@ Appendix B Cost Considerations
The availability of open LLMs such as Llama3.1 or Qwen2.5 present serious competition to commercial services like the hosted GPT-4o (-mini) models. However, the price of of $1.25$ USD for 1M input tokens of OpenAI’s batch API may be considered as relatively low compared, due to a rapid fall in prices compared to legacy models such as GPT-4. Hence, the fundamental choice between commercial or open LLMs depends not only on the expected scale of the amount of data to process or the size of the datasets, but also on data protection considerations, that are crucial for sensitive legal data.

### v2

@@ 6.3 Limitations and Outlook
We see several important directions for future work. First, open questions remain as to potential improvements when combining the most advanced architectures and training methodologies, such as DeBERTaV3, with a suitable language- and domain-specific training corpora, such as those used by RoBERTalex or BETO. Analogously to domain-specific pre-training, [24] propose the Lawma model by fine-tuning Llama-3 8b on 260 legal tasks, effectively emulating pre-training. Their Lawma outperforms GPT-4 in the zero- and three-shot settings. We consider both Spanish legalese pre-training and task-specific fine-tuning of LLMs with more than a billion parameters an interesting avenue for future work, as [24] report increased performance from only a few hundred samples. The findings of [9] on the limited reasoning capabilities of LLMs even necessitate enhancements of today’s generative LLMs beyond zero-shot learning.
@@ 7 Conclusion
A growing number of products and services are offered through online platforms. However, their Terms of Service remain enigmatic to consumers. Even if they read them, there is no alternative but to accept them. This rational apathy and information asymmetry present in modern consumer contracts may result in uncontested but potentially abusive clauses and their adverse effects on consumers. It follows that there is a growing interest in the potential of automatic contract analysis.
We present a novel annotation scheme and a substantial dataset comprising 50 online Terms of Service. The corpus encompasses two tasks for three categories of potentially abusive clauses. Our extensive evaluation of a diverse set of transformer-based methods reveals that multiple factors impact the detection and classification rates, including language- and domain-specific pre-training, the number of samples in few-shot prompts, and architectural considerations. Potentially abusive clauses are detected with macro-F1 scores ranging from between $79\%$ to $89\%$ and micro-F1 scores of $95\%$ to $96\%$. The highest classification rates range from between $60\%$ to $70\%$ macro-F1 score and $64\%$ to $80\%$ micro-F1 score. This underscores the wide range of difficulties inherent in the novel datasets.
This work introduces the first Spanish-language classification dataset of legal clauses. Moreover, it applies law from outside a major jurisdiction, specifically the Chilean legal code. We also present the first comprehensive study on Spanish-language and multilingual transformer-based models on legal classification tasks in Spanish. This paper serves as the foundation for substantial future research and the implementation of its methodology to benefit Chilean and Latin American consumers.
@@ Appendix B Cost Considerations
The availability of open LLMs such as Llama3.1 or Qwen2.5 presents serious competition to commercial services like the hosted GPT-4o (-mini) models. However, the price of $1.25$ USD for 1M input tokens of OpenAI’s batch API may be considered relatively low, due to a rapid fall in prices compared to legacy models such as GPT-4. Hence, the fundamental choice between commercial or open LLMs depends not only on the expected scale of the amount of data to process or the size of the datasets, but also on data protection considerations, that are crucial for sensitive legal data.